OpenStudy (anonymous):

z^2=7-6i; solve for z

OpenStudy (wmj259):

assume z is y and i is x

OpenStudy (wmj259):

\[y ^{2}=7-6x\]

OpenStudy (wmj259):

\[\sqrt{y ^{2}}=\sqrt{7-6x}\]

OpenStudy (wmj259):

\[y=\sqrt{7-6x}\]

OpenStudy (wmj259):

Just switch y for z and x for i

TheSmartOne (thesmartone):

So after doing what @wmj259 did you get \[z=\sqrt{7-6i}\]

OpenStudy (paxpolaris):

we can do better than that, no?

TheSmartOne (thesmartone):

I don't think so...

geerky42 (geerky42):

Why is it necessary to replace z and i to y and x respectively?

OpenStudy (freckles):

\[\text{ well you can write } 7-6i \text{ in } re^{i \theta} \text{ form }\]\[z^2=r (\cos(\theta)+i \sin(\theta))=r(\cos( \theta+ 2 \pi n)+i \sin(\theta+2 \pi n)) =re^{i (\theta+2 \pi n)} \\ z=r^\frac{1}{2}e^{\frac{1}{2}i (\theta+ 2 \pi n)}\] And we only need to two z values since we are finding square root so n=0,1 in this case

OpenStudy (freckles):

All you have to do is find r and thet

OpenStudy (paxpolaris):

 \[y^2=7-6i = \sqrt{85}e^{i \theta}\] i guess for the exact form of y we'd have to leave the expression as\[\large y = \color{red}\pm \sqrt{7-6i}\]

OpenStudy (wmj259):

@geerky42, you don't have the change the variables, but its easier to see it.

OpenStudy (wmj259):

that is true that when you take the square root you have to include the plus or minus

OpenStudy (paxpolaris):

the approximate form would be:\[y \approx \pm2.8478\mp1.0535 i\]